OpenStudy (anonymous):

I need help to solve this question? A large department store chain is considering opening a new store in a town of 15,000 people. Before making the decision, a market survey is conducted. Of 200 persons interviewed, 165 indicated they would patronize the new store. Find a 95% confidence interval on the proportion p of people who will patronize the store. Use this to give a 95% confidence interval on the number of people who will patronize the store.

OpenStudy (amistre64):

do you know the zscore of 95%?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

\[\hat p\pm Z_{95}~\sqrt{\frac{pq}{n}}\]

OpenStudy (amistre64):

you have to have a way to determine the zscore for 95% ... what methods are available to you?

OpenStudy (anonymous):

The range of the confidence interval is defined by the sample statistic + margin of error. And the uncertainty is denoted by the confidence level.

OpenStudy (amistre64):

\(\hat p\) is the fraction of people that liked it: 165 out of 200 the Z is the associated zscore for 95% about the mean and the rest is really amounts to the square root of: 165 times (200-165), divided by 200^3

OpenStudy (amistre64):

lets assume for the moment that a zscore of 95% is say ... 1.96

OpenStudy (amistre64):

\[\frac{165}{200}\pm1.96\sqrt{\frac{165*35}{200^3}}\]represents our confidence interval

OpenStudy (anonymous):

you get the value of 1.96 from the table

OpenStudy (anonymous):

thank you very much

OpenStudy (amistre64):

youre welcome